A car of mass 1000 kg moving witha velocity of 90 km/h collides with a tree and comes to stop in 10 s. What is the change in momentum of the car? What is the force applied by the tree to stop the car?
1
Expert's answer
2019-10-21T08:53:23-0400

The change in momentum of the car


"\Delta p=m|v_f-v_i|"

"=1000\: \rm kg\times |0\: m/s-90\: \frac{1000\: m}{3600\: s}|=25000\: kg\cdot m/s"

The force applied by the tree to stop the car is equal


"F=\frac{\Delta p}{\Delta t}=\rm\frac{25000\: kg\cdot m/s}{10\: s}=2500\: N"